OpenStudy (anonymous):

Write a formula for this sequence: 50, 83, 116, 149,... 33n + 50 33n + 17 149 - 33n Could you at leasttell me how to do this?? please and thank you!!

OpenStudy (anonymous):

First see the distance between 50 and 83. So subtract. 83 - 50 = 33 Then add 33 to 83 to see if that is the correct pattern. 83 + 33 = 116.

OpenStudy (anonymous):

yhow to find the common difference but how do i find the formula?

OpenStudy (anonymous):

I would pick B as the correct answer!

OpenStudy (cwrw238):

yes because when n = 1 first term = 33(1) + 17 = 50 and third term = 33(3) + 17 = 116